OpenStudy (jadedry):

The deadline for this question has passed, so I feel as if it is safe to post here:

OpenStudy (jadedry):

The rate constants for a reaction are found to be k(110 C) = 7.0 x 10 -4 s-1 k(165 C) = 41.9 x 10 -4 s-1 Determine the rate constant in units of s-1 at 269 C For starters, I have /just/ started 11th grade so I know very little about calculus, I am familiar with logarithms though. I know I have to answer this question using the Rate of Reaction Equation , but I don't know what the symbols in the equation mean, or how to apply it to this question. Any hints? Thanks in advance! c:

OpenStudy (jadedry):

@Rushwr Hope you don't mind the ping! ;u;

OpenStudy (anonymous):

what you need to do is use the arrehnius parameter: \[k=Ae ^{-\frac{ E _{a} }{ RT }}\]

OpenStudy (anonymous):

perhaps take the logarithm on both sides to get \[\ln(k)=\ln(A)-\left( \frac{ E }{ R } \right)\left( \frac{ 1 }{ T } \right)\]

OpenStudy (anonymous):

we know R right, 8.314, etc...universal gas constant

OpenStudy (anonymous):

we also know two rate constants at its specific temperature. Therefore, we can make a system of two linear equations

OpenStudy (anonymous):

to solve for the pre-exponential factor and the activation energy

OpenStudy (anonymous):

then we just simply use the general rate equation and just substitute the known pre-exponential factor and the activation energy values as well as the temperature at which we are asked to find the rate constant.

OpenStudy (jadedry):

Yeah, Boltzman Constant I believe? I see where you're going with this. So this is what it should look like with certain values plugged in before I solve the linear equations? \[\ln(7*10^{-4}) = \ln(A) - \frac{ E }{ 8.314} * \frac{ 1 }{ 110}\] \[\ln(41.9*10^{-4}) = \ln(A) - \frac{ E }{ 8.314} * \frac{ 1 }{ 165}\]

OpenStudy (anonymous):

make sure your temperature is in kelvin

OpenStudy (jadedry):

\[\ln(7*10^{-4}) = \ln(A) - \frac{ E }{ 3185.5091} \] (I converted C to K) \[\ln(41*10^{-4}) = \ln(A) - \frac{ E }{ 3642.7791} \] So subtracting the first from the second: \[\ln\frac{ 41*10^{-4} }{ 7*10^{-4} } = \frac{ E }{ 3642.78 } - \frac{ E }{ 3185.5091 }\]

OpenStudy (jadedry):

Ah sorry, was a bit behind you, let me catch up.

OpenStudy (anonymous):

oops, my equations are slightly incorrect. the lHS of each equation should have ln(..)

OpenStudy (anonymous):

i'll re type

OpenStudy (jadedry):

My \[\frac{ E }{ 3642.78}\] looks different to yours? perhaps I did something wrong? (I checked my first E as well, it's the same)

OpenStudy (anonymous):

yeah my bad

OpenStudy (anonymous):

i'll fix it again

OpenStudy (jadedry):

I feel your pain haha.

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

\[\ln(7\times10^{-4})=lnA-3.14\times10^{-4}E\] \[\ln(41.9\times10^{-4})=lnA-2.75\times10^{-4}E\]

OpenStudy (anonymous):

thats better

OpenStudy (anonymous):

so i'm going to do it quick, but what you should get is ln(A)=7.1423 E=45881

OpenStudy (anonymous):

hence we know have a general solution for such rate constant at any temperature. \[\ln(k)=7.1423-\left( \frac{ 45881.43}{ 8.314 } \right)\left( \frac{ 1 }{ T } \right)\] Simply just sub T=269 into the equation to get a value for ln(k) \[\ln(k)=7.1423-\left( \frac{ 45881.43}{ 8.314 } \right)\left( \frac{ 1 }{ 269+273 } \right)\] Then evaluating \[\ln(k)=-3.03957...\] Taking exponential to isolate k, rate constant at 269 degrees celcius is: \[k=0.048 {\space} s ^{-1}\]

OpenStudy (jadedry):

Okay After that: \[\ln \frac{ 41*10^{-4} }{ 7*10^{-4} } =0.39*10 ^{-4} E\] And then you take the value of E, stick it in there and find A, Gotcha. Then you plug them into the first equation you gave me, the Arrhenius parameter? Gotcha, it's pretty clear now. I was staring at the equation for general molecules, super confusing.

OpenStudy (jadedry):

Thanks a lot! Your detailed explanations were just what I was looking for. c:

OpenStudy (anonymous):

remember, R is not the baltzman constant. Thats a different variable. Here, R is the universal gas constant

OpenStudy (jadedry):

Yes, my bad. But essentially it's the same value of Boltzman constant but with different units? Right? (hence a different number)

OpenStudy (anonymous):

remember when you deal with rate constants always think KINETICS. the Arrehnius equation is so fundamental for reaction kinetics

OpenStudy (anonymous):

yeah pretty much

OpenStudy (anonymous):

i completed a 3rd year uni kinetic and reactor design course, so anything reaction rates, i'll boss

OpenStudy (jadedry):

Alright, I understand. I wasn't aware of that equation before. Just want to confirm, E was activation energy and A was the pre-exponential constant? Impressive! If I have trouble with this stuff in future I hope you won't mind if I give you a ping haha.

OpenStudy (anonymous):

yep that is correct! so you never came across that equation? hmm

OpenStudy (jadedry):

Well, I probably /did/ but in such a way that it wasn't highlighted. I'm doing the EDX MIT Solid Chem course, and I'm struggling because I don't even know some of the math I should.

OpenStudy (jadedry):

Either way I know know, so if I some something in a similar vein I'll kick its a**

OpenStudy (jadedry):

*know now

OpenStudy (anonymous):

haha awesome! good luck with it! i'm off to bed. catchya

OpenStudy (jadedry):

Good night! Thanks again!